OpenStudy (anonymous):

Find dy/dx in terms of x and y if x^2 y - x - 4 y - 10 =0. dy/dx =

OpenStudy (redheadangel):

wow this is like college level stuff, maybe @Here_to_Help15 can help! (lol i always reference him)

OpenStudy (anonymous):

\[(x^2'y+xy')-1-4y'=0\] I thought you'd start like this but then I ended up getting the wrong answer.

OpenStudy (cwrw238):

x^2 y - x - 4 y - 10 =0 (x^2 * dy/dx + 2xy) - 1 - 4 dy/dx = 0

OpenStudy (cwrw238):

now make dy/dx the subject

OpenStudy (cwrw238):

ok - can you continue?